# Attention-Aware Visualization: Tracking and Responding to User Perception Over Time

### v1

@@ 1 Introduction
In this paper, we introduce the concept of attention-aware visualizations (AAVs): contextually aware visual representations that track user attention on their surface over time and modify their visual attributes accordingly. Conceptually, this approach consists of three logical steps:
@@
    Recording attention over time as the user’s perception moves over the visualization. This also includes modeling the user’s short-term memory as attention smoothly decaying over time.
@@
The contributions of this paper are as follows: (1) the concept of attention-aware visualization as a context-aware [39] method for creating visualizations responsive to implicit gaze and not just explicit interaction; (2) a library implementing attention-awareness for 2D visualizations based on numeric integration; (3) an implementation of attention-awareness and occlusion [16] for 3D visualizations using the stencil buffer; and (4) results from a user study investigating the utility of the approach for both 2D (eye-tracker) and 3D settings (XR HMD). [^1] Research materials available on OSF: https://osf.io/8mfhp/
@@ 2.2 Attention in Visualization
The human perceptual and cognitive systems generate a wealth of information in real time, all of which cannot be processed simultaneously. Attention is the selective concentration of mental awareness on a subset of this information, excluding the remaining information [44]. Early work on cognitive processes in attention noted that for highly complex attentive processes—such as visual search—to be possible, they must be organized in both low-level parallel—so called pre-attentive—and complex serial—attentive—processes [23, 31]. While subsequent work has criticized the existence of a clear division between pre-attentiveness and attentiveness [52], it remains a useful construct for understanding operational aspects of human perception and attention.
@@ 2.4 Gaze as Input
The last decade has seen significant advancements in gaze-supported multimodal interactions, particularly for exploring large datasets and improving target acquisition.
Stellmach et al. [43] and Stellmach and Dachselt [42] explored gaze-supported interactions for large image collection exploration and target acquisition, respectively.
Turner et al. [46], Pfeuffer et al. [34, 35], and Velloso et al. [47] further investigated the integration of gaze with touch and mid-air gestures, demonstrating the potential for seamless and efficient interactions across devices and interfaces.
Pfeuffer et al.’s work on combining gaze with pinch interactions in virtual reality [36] represents the latest evolution of this research direction, highlighting the growing sophistication and breadth of gaze-supported interactions. [^2] Incidentally, gaze and pinch has also been adopted by the Apple Vision Pro, making it the first commercial device using dedicated gaze interaction.
This is also the research most closely related to our attention-aware visualizations, but our approach is less concerned with direct gaze input to control an interface, and more with retrospective analysis of gaze patterns over time.
@@ 3 Design: Attention-Aware Visualization
The intellectual contribution proposed in this work is the concept of attention-aware visualizations (AAV): data visualizations that track the viewer’s attention on their surface over time and react accordingly. We see several benefits of this approach, from creating a more responsive and engaging visual environment to showing data coverage, illuminating missed outliers, or even predicting attention patterns. We can summarize our approach using the following research questions:
@@ 3.3 Revisualizing Attention
How do we actually use the accumulated attention metric to change the visualization (RQ3)? One option is to not change the visualization at all, but to merely use the data for evaluation purposes where the viewers themselves do not ever see it. However, in this paper, we are interested in investigating how this data can be utilized by the visualization itself. We call this revisualizing the attention on the original visualization.
@@
    Border: Use view border to visualize attention; and
@@ 4.2 Triggering Mechanisms
As noted in Section 3.3, we identify three strategies for triggering the revisualization of the visualization: Always-on, Explicit triggering, and Implicit triggering (Figure 1). Always-on is constantly active. Depending on the revisualization approach, this behavior may lead to the revisualization itself affecting the user’s attention.
@@ 4.3 Revisualizing Attention
[Glaze overlay.] Refer to caption [Decorative frame border.] Refer to caption [Minimap on the border.] Refer to caption
@@ 5.2 Revisualization
Furthermore, AAV2D supports several revisualization approaches depending on the presentation method. For both the glaze overlay and border minimap, the approach can use a heatmap, contour plot, or (Figure 3); in the case of the overlay, this revisualization is shown on top of the element, whereas for the minimap, it is shown on the miniature. For the frame border, we support bar charts, area charts, and heatmaps that show the short-term or cumulative attention on the plot for a specific $x$ or $y$ coordinate; see Figure 4 for examples.
@@ 5.3 Implementation Notes
We implemented AAV2D using vanilla JavaScript without the use of specialized libraries or bindings. D3333https://d3js.org [10] was used for composing Scalable Vector Graphics (SVG) of contour and heatmaps. The mesh implementation uses CSS image filters to perform some of its visual transformations on the underlying HTML element, such as blur, saturation, and hue-rotation. All input events are intercepted by AAV2D event handlers, but are then propagated to the underlying element, meaning that it can be used even for interactive visualizations and interfaces.
@@ 6.1 Attention Capture
The second strategy only apply attention to elements in the center of the screen. This strategy is specifically designed for immersive environments using XR headsets that do not include eyetracking, such as the Meta Quest 3. In such immersive environments, the user’s head orientation thus acts as a proxy for their attention.
@@ 6.3 Implementation Notes
Our 3D visualization is implemented as a WebXR application using the Three.js library. [^4] https://threejs.org Our visualizations are bespoke; i.e., we do not use a specific 3D visualization library. Attention on objects is achieved through GPU-picking, where each frame is rendered first in a picking scene, and then in the normal scene, which is shown on screen. Each triangle of every object’s geometry is assigned a unique RGB color in the picking scene. The eight red color bits are used to identify the object, while the 16 green and blue color bits are used to identify faces in the object geometry. This way allows us to know which specific part of the object to add attention to. Figure 6 illustrates how an object—in this case a low-polygon sphere—is colored in the picking scene.
Figure 6: GPU picking based on IDs. Each face of every object is assigned a color in the picking scene, based on the ID of the object and the ID of the face of that object. The red RGB colorspace is reserved for object IDs, while the green and blue RGB colorspaces are used for faces. The colors in this image are exaggerated to make them distinguishable.
The current implementation has a few limitations. Firstly, GPU picking is in itself an expensive operation. To reduce the impacts this operation has on performance, we do not perform this operation at every frame, but rather at a specified interval of 100 milliseconds. Secondly, the use of colors limits the number, and complexities, of objects in the scene, and thus the size of the dataset to be visualized. This has not been a problem for the datasets we have used.
@@ 7 Evaluation Study
Figure 7: Experimental interface. 2D and 3D user study setups.
@@ 7.2 Participants
We recruited 12 participants (9 male, 3 female) from a convenience population (students at two different universities). The participants were screened to be experienced computer users and knowledgeable about data visualization. Their age ranged from 24 to 52 (mean 30), all were enrolled in a computer science program, and all had normal or corrected-to-normal vision with no color vision deficiencies.
@@ 7.4 Procedure
Participants first provided informed consent at the beginning of their evaluation session. They filled out a demographic form and were given a brief introduction. They were then asked to complete three task blocks in random order. Each task block consisted of a visualization and a list of four open-ended questions. The visualizations are shown in Figure 8. Participants were assigned a random combination of attention presentation and triggering mechanism. Table 2 gives the experimental conditions, three for 2D and 3D each. Participants were given unlimited time to find answers to questions. Their performance was not timed and correctness was not measured. Participants were encouraged to verbalize their thinking, but did not follow a strict think-aloud protocol.
@@ 7.5.1 Observations
Participants were all able to complete tasks. We made no specific direct observations about task performance during either the 2D or 3D study.
We observed that the Pupil Labs Neon eye-tracker is not completely accurate. For some participants it measured slightly below their actual gaze, and for others slightly above, which likely skewed the attention tracking. This will slightly offset the measured attention on the visualization, which leads to incorrect attention capture and corresponding inaccurate revisualizations. However, we did not observe this being an issue for our study, as the tasks were not dependant on revisualization accuracy, and the participants did not voice any accuracy concerns.
@@ 7.5.2 Subjective Ratings
The results for the SUS questionnaire were overall just above average [6], with the average score across participants of 70 (SD = 15). Per condition scores were all around 65 (below average SUS score), with one notable exception, the 3D visualization with explicit triggering, which scored 86 (excellent SUS score). When asked to rate the usefulness of the revisualizations, the responses varied significantly, with slightly more positive than negative responses. Figure 9 illustrates the reported usefulness of all the revisualizations. These dispersed responses aligns with the SUS questionnaire results.
When asked to rank the triggering and revisualization techniques, the participants using the 3D implementation had a clear preference for the explicit triggering condition and disliked the always-on heatmap. They preferred having control over when the revisualization was triggered, noting that the constant recoloring in the always-on condition was highly distracting, particularly because it overrode the colors of the visualization itself. This preference for explicit triggering is supported by the SUS score. The implicit triggering conditions were considered hard to understand, with one participant referring to it as “mysterious.” At the same time, other participants considered implicit triggering to be useful for certain tasks by helping to direct their attention.
@@ 7.5.3 Participant Comments and Feedback
Revisualizations were predominantly used by participants to refocus and redirect their attention, often illuminating areas within the visual space that had either been overlooked or required further exploration. This utility of revisualizations bridged both 2D and 3D contexts, enhancing users’ exploration thoroughness by guiding them towards both explored and unexplored areas (P2, P3, P4, P5 in 2D; P1, P3, P5 in 3D). A participant (2D, P2) highlighted this dual functionality, noting, “It was more that it guided me to where I had been looking and which, of course, did tell me where I haven’t been looking.” Similarly, the stimulation of curiosity, especially when participants were seeking specific details or understanding the breadth of visualization techniques, was a recurring theme (P5, P2 in 2D; P2, P3 in 3D). Some participants in 3D shared that they “treated it like a game”(3D, P4) and “found (themselves) wanting to go inside the visualization and see it from inside,” (3D, P2), underscoring the depth of engagement revisualizations can foster.
@@
Participant feedback also pointed to the highly subjective nature of revisualization preferences and the critical role of task suitability in the effectiveness of these tools. This was evident in the 2D context, as some participants gravitated towards specific revisualization techniques such as heatmaps or contour maps based on their orientation towards precision and their detail-oriented utility, others valued minimaps for providing a holistic perspective of the data (P3, P2, P4, P5, P6 in 2D; P3, P5, P6 in 3D). “I did use the heatmap and the contour map more because I found them easier to understand how they can help me,” noted a participant (2D, P3), exemplifying the varied preferences among users. This diversity underscores the need for adaptable revisualization techniques that can cater to the preferences of individual users and specific analytical tasks.
@@
Participants increasingly engaged with the revisualization techniques as the study progressed. Initially drawn to these tools by their novelty, participants experienced a rise in their comfort and efficiency in using them over time (P4 in 2D; P1, P2, P3, P5 in 3D). One participant in the 2D setting remarked, “At first, I was really paying attention because it was new… but after getting familiar, it feels routine and I hardly give it a second thought” (P4 in 2D). Similarly, in the 3D setting, another participant noted, “Indeed, as I became more accustomed to the effects, I found my attention more effectively directed,” highlighting the dynamic process of their interaction with the tools (3D, P1). This suggests that over time, users are likely to not just get used to the revisualization features but also incorporate them more fluidly into their analytical processes.
@@ 8.2 Design Implications
Furthermore, while we left collaboration out of our treatment in this paper, the step to collaborative and educational settings is not far. In environments where multiple users interact with the same visualization, AAVs could dynamically adjust to accommodate the focus and needs of different users, enhancing collaborative data analysis and decision-making. In educational contexts, AAVs could tailor the presentation of information to the learner’s pace and points of interest, providing a more engaging and effective learning experience. Realizing these futures means developing sophisticated algorithms capable of managing the delicate balance between being informative and non-disruptive, ensuring that AAVs enrich the user experience by making data visualizations not just tools for viewing but also guiding exploration.

### v2

@@ Abstract
A teaser image for Attention Aware Visualization Paper. The image is divided into three main sections: 1. Middle Section: Two figures: (a) Left figure with VR headset labeled "Data Aware 3D" with Meta Quest 3 surrounded by 3 icons. Icons represent orientation, rotation, and location. (b) Right figure with eye-tracking device labeled "Data Agnostic 2D" with Pupil Labs Neon Eyetracker surrounded by 3 icons. Icons represent gaze, pointer, and keys. 2. Left Section shows Data Aware 3D Implementation: (a) Explicit Heatmap: Shows 3D bar graphs with heatmaps. (b) Implicit Desaturation: Displays scatter plots for "Sepal Width" with gradual desaturation. (c) Always On Heatmap: Shows 3D surface plots with heatmaps for "Latitude" and "Longitude." 3. Right Section shows Data Agnostic 2D Implementation in two parts: Part 1. IllustratedPicture Framing Metaphor. Part 2. Three types of frames: (a) Bar Frame: Bar graph with maximum cumulative attention. (b) Area Frame: Area plot showing attention distribution. (c) Heat Frame: Continuous heatmap gradient.
@@ 1 Introduction
We introduce the concept of attention-aware visualizations (AAVs): contextually aware visual representations that track user attention on their surface over time and modify their visual attributes accordingly. Conceptually, this approach consists of three steps (Attention-Aware Visualization: Tracking and Responding to User Perception Over Time):
@@
    Recording attention over time as the user’s perception (i.e. gaze) moves over the visualization over time. This also includes modeling the user’s short-term memory as attention smoothly decaying.
@@
While there are numerous potential approaches to designing AAVs, this paper focuses on the fundamental idea of showing users their own attention on a visualization. Our goal is to encourage users to explore areas of a visualization they may not have previously investigated as well as to reflect on the areas they have, thus enhancing their overall understanding and interaction with the data.
The contributions of this paper are as follows: (1) the concept of attention-aware visualization as a context-aware [41] method for creating visualizations responsive to implicit gaze and not just explicit interaction; (2) a library implementing attention-awareness for 2D visualizations based on numeric integration; (3) an implementation of attention-awareness and occlusion [16] for 3D visualizations using the stencil buffer; and (4) results from a qualitative evaluation investigating the approach for both 2D (eye-tracker) and 3D settings (XR HMD). Refer to Supplemental Materials for the supplementary materials.
@@ 2.2 Attention in Visualization
The human perceptual and cognitive systems generate a wealth of information in real time, all of which cannot be processed simultaneously. Attention is the selective concentration of mental awareness on a subset of this information, excluding the remaining information [46]. Early work on cognitive processes in attention noted that for highly complex attentive processes—such as visual search—to be possible, they must be organized in both low-level parallel—so called pre-attentive—and complex serial—attentive—processes [23, 33]. While subsequent work has criticized the existence of a clear division between pre-attentiveness and attentiveness [54], it remains a useful construct for understanding operational aspects of human perception and attention, e.g., for visual search optimization [29].
@@ 2.4 Gaze as Input
Over the past decade, there has been significant progress in gaze-supported multimodal interactions, particularly for exploring large datasets and improving target acquisition. Stellmach et al. [45] and Stellmach and Dachselt [44] explored gaze-supported interactions for large image collection exploration and target acquisition, respectively. Turner et al. [48], Pfeuffer et al. [36, 37], and Velloso et al. [49] further investigated the integration of gaze with touch and mid-air gestures, demonstrating the potential for seamless and efficient interactions across devices and interfaces. Pfeuffer et al.’s work on combining gaze with pinch interactions in virtual reality [38] represents the latest evolution of this research direction, highlighting the growing sophistication and breadth of gaze-supported interactions. [^1] Incidentally, gaze and pinch has also been adopted by the Apple Vision Pro, making it the first commercial device using dedicated gaze interaction. Although this research is closely related to AAVs, our approach focuses less on utilizing direct gaze as an input to control an interface, as explored in [28], and more on the retrospective analysis of gaze patterns over time.
@@ 3 Design: Attention-Aware Visualization
The intellectual contribution proposed in this work is the concept of attention-aware visualizations (AAV): data visualizations that track the viewer’s attention on their surface over time and react accordingly. We see several benefits to this approach, ranging from creating a more responsive and engaging visual environment, to showing data coverage, illuminating missed outliers, and even predicting attention patterns. Although numerous possibilities exist for how a visualization could respond to a viewer’s attention, this paper focuses solely on displaying the viewer’s attention on the visualization. We can summarize our approach using the following research questions:
@@ 3.3 Revisualizing Attention
How do we actually use the accumulated attention metric to change the visualization (RQ3)? One option is to not change the visualization at all, but to merely use the data for evaluation purposes where the viewers themselves do not ever see it. However, in this paper, we are interested in investigating how this data can be utilized by the visualization itself, e.g. to encourage exploration, or facilitate introspection. We call this revisualizing the attention on the original visualization.
@@
    Border: Use view borders to visualize attention; and
@@ 4.2 Triggering Mechanisms
As noted in Section 3.3, we identify three strategies for triggering the revisualization of the visualization: Always-on, Explicit triggering, and Implicit triggering (Fig. 1). Always-on is constantly active. Depending on the revisualization approach, this behavior may lead to the revisualization itself affecting the user’s attention.
@@ 4.3 Revisualizing Attention
The use of color to encode the revisualization of course means that color cannot be used to simultaneously encode dimensions of the data, which is not desired. However for immersive environments, using e.g. scale to denote attention would skew the perception of the data. For this paper we consider it a necessary compromise, and using an explicit triggering approach still allows for encoding data using color.
[Glaze overlay.] Refer to caption [Minimap on the border.] Refer to caption [Decorative frame border.] Refer to caption
@@ 5.2 Revisualization
Furthermore, AAV2D supports several revisualization approaches depending on the presentation method. For both the glaze overlay and border minimap, the approach can use a heatmap, contour plot, or mesh (Fig. 3); in the case of the overlay, this revisualization is shown on top of the element, whereas for the minimap, it is shown on the miniature. For the frame border, we support bar charts, area charts, and heatmaps that show the short-term or cumulative attention on the plot for a specific $x$ or $y$ coordinate; see Fig. 4 for examples.
@@ 5.3 Implementation Notes
We implemented AAV2D using vanilla JavaScript without the use of specialized libraries or bindings. D3222https://d3js.org [10] was used for composing Scalable Vector Graphics (SVG) of contour and heatmaps. The mesh implementation uses CSS image filters to perform some of its visual transformations on the underlying HTML element, such as blur, saturation, and hue-rotation. All input events are intercepted by AAV2D event handlers, but are then propagated to the underlying element, meaning that it can be used even for interactive visualizations and interfaces.
@@ 6.1 Attention Capture
The second strategy only apply attention to elements in the center of the screen. This strategy is specifically designed for immersive environments using XR headsets that do not include eyetracking, such as the Meta Quest 3 (which was the target implementation platform for AAV3D). In such immersive environments, the user’s head orientation thus acts as a proxy for their attention.
@@ 6.3 Implementation Notes
Our 3D visualization implements bespoke visualizations in a WebXR application using the Three.js library. [^3] https://threejs.org Attention is tracked through GPU color-picking on a per face basis; i.e., each triangle of every object’s geometry is assigned a unique RGB color in the picking scene. The red RGB colorspace is reserved for object IDs, while the green and blue RGB colorspaces are used for faces. A benefit of using GPU-picking is that it is independent of the data geometry; this means we can even use custom 3D models. With this approach we also do not need to be concerned with the layout of the 3D scene, and we can more easily get objects from an area compared to using a raycaster.
The current implementation has a few limitations. Firstly, GPU picking is in itself an expensive operation. To improve performance, we only perform color-picking every 100 milliseconds, rather than every frame. Secondly, the use of colors limits the number, and complexities, of objects in the scene, and thus the size of the dataset to be visualized. This has not been a problem for the datasets we have used.
@@ 7.1 Study Design Rationale
Figure 6: Experimental interface. 2D and 3D user study setups.
@@ 7.2 Participants
We recruited 12 participants (9 male, 3 female) from a convenience population (students at two different universities). The participants were screened to be experienced computer users and knowledgeable about reading and interpreting data visualizations through a self assessment questionaire. Their age ranged from 24 to 52 (mean 30), all were enrolled in a computer science program, and all had normal or corrected-to-normal vision with no color vision deficiencies.
@@ 7.4 Procedure
Participants first provided informed consent at the beginning of their evaluation session. They filled out a demographic form and were given a brief introduction. They were then asked to complete three task blocks in random order. Each task block consisted of a visualization and a list of four open-ended questions. For the 2D tasks, participants engaged with 3 visualizations: Charles Minard’s Napoleon March (Fig. 7), John Snow’s Cholera Map (Fig. 7), and Florence Nightingale’s Rose Chart (Fig. 7). They were asked questions warranting exploration, such as identifying river crossings by Napoleon’s army, counting the number of pumps contaminated by cholera, and comparing mortality rates from diseases across different years. For the 3D tasks, participants worked with a scatterplot of R.A. Fisher’s Iris data set (Fig. 7), a barchart of Quinlan’s Auto MPG dataset from UC Irvine (Fig. 7), and an elevation map of Mount St. Bruno (Fig. 7). They were asked to analyze car cylinder counts over time, identify the largest flower sepals, and compare elevations on the terrain map. Detailed questions for these tasks can be found in the supplemental materials (see Supplemental Materials). Participants were assigned a random combination of attention presentation and triggering mechanism. Table 2 gives the experimental conditions, three for 2D and 3D each. Participants were given unlimited time to find answers to questions. Their performance was not timed and correctness was not measured. Participants were encouraged to verbalize their thinking, but did not follow a strict think-aloud protocol.
@@ 7.5.1 Observations
All participants were able to complete all the tasks. We made no specific direct observations regarding task performance during either the 2D or 3D study. We observed that the Pupil Labs Neon eye-tracker used for the 2D study was not completely accurate. For some participants it measured slightly below their actual gaze, and for others slightly above, which likely skewed the attention tracking. Although this slightly offset measured attention on the visualization potentially led to inaccurate revisualizations, we did not observe this being an issue for our study, given that the tasks were not dependent on revisualization accuracy, and the participants did not voice any accuracy concerns.
@@ 7.5.2 Subjective Ratings
The results for the SUS questionnaire were overall just above average [6], with the average score across participants of 70 (SD = 15). Per condition scores were all around 65 (below average SUS score), with one notable exception, the 3D visualization with explicit triggering, which scored 86 (excellent SUS score). When asked to rate the usefulness of the revisualizations, the responses varied significantly, with slightly more positive than negative responses. Fig. 8 illustrates the reported usefulness of all the revisualizations. These dispersed responses aligns with the SUS questionnaire results.
When asked to rank the triggering and revisualization techniques, participants using the 3D implementation had a clear preference for the explicit triggering condition and disliked the always-on heatmap. They preferred having control over when the revisualization was triggered, noting that the constant recoloring in the always-on condition was highly distracting, particularly because it overrode the colors of the visualization itself. This preference for explicit triggering is supported by the SUS score. The implicit triggering conditions were considered hard to understand, with one participant referring to it as “mysterious.” At the same time, other participants considered implicit triggering useful for certain tasks by helping to direct their attention.
@@ 7.5.3 Participant Comments and Feedback
Revisualizations were predominantly used by participants to refocus and redirect their attention, often illuminating areas within the visual space that had either been overlooked or required further exploration. This utility of revisualizations bridged both 2D and 3D contexts, enhancing users’ exploration thoroughness by guiding them towards both explored and unexplored areas (P2–P5 in 2D; P1, P3, P5 in 3D). A participant (2D, P2) highlighted this dual functionality, noting, “It was more that it guided me to where I had been looking and which, of course, did tell me where I haven’t been looking.” Similarly, the stimulation of curiosity, especially when participants were seeking specific details or understanding the breadth of visualization techniques, was a recurring theme (P2, P5 in 2D; P2, P3 in 3D). Some participants in 3D shared that they “treated it like a game”(3D, P4) and “found (themselves) wanting to go inside the visualization and see it from inside,” (3D, P2), underscoring the depth of engagement revisualizations can foster.
@@
Participant feedback also pointed to the highly subjective nature of revisualization preferences and the critical role of task suitability in the effectiveness of these tools. This was evident in the 2D context, as some participants gravitated towards specific revisualization techniques such as heatmaps or contour maps based on their orientation towards precision and their detail-oriented utility, others valued minimaps for providing a holistic perspective of the data (P2–P6 in 2D; P3, P5, P6 in 3D) “I did use the heatmap and the contour map more because I found them easier to understand how they can help me,” noted a participant (2D, P3), exemplifying the varied preferences among users. This diversity underscores the need for adaptable revisualization techniques that can cater to the preferences of individual users and specific analytical tasks.
@@
Participants increasingly engaged with the revisualization techniques as the study progressed. Initially drawn to these tools by their novelty, participants experienced a rise in their comfort and efficiency in using them over time (P4 in 2D; P1–P3, P5 in 3D). One participant in the 2D setting remarked, “At first, I was really paying attention because it was new… but after getting familiar, it feels routine and I hardly give it a second thought” (P4 in 2D). Similarly, in the 3D setting, another participant noted, “Indeed, as I became more accustomed to the effects, I found my attention more effectively directed,” highlighting the dynamic process of their interaction with the tools (3D, P1). This suggests that over time, users are likely to not just get used to the revisualization features but also incorporate them more fluidly into their analytical processes.
@@ 8.2 Design Implications
Even though we have only implemented and evaluated two of the settings that AAV could be useful in, we feel that AAV has potential in other settings as well, such as attention-aware 2D visualizations in immersive environments, or 3D visualizations viewed in a non-immersive environment, i.e., on a screen. Non-immersive 3D visualizations are arguably more common than their immersive counterparts, which is why our approach in AAV3D is designed to also work for such settings. We chose to focus on an immersive environment in this paper because we believe that AAVs can provide especially great value for such settings. When data is distributed all around the user, the cognitive load of keeping track of the data is high. In such scenarios, AAVs could be used to reduce the cognitive load, by pointing to unseen areas, or by reminding the user if they have already analyzed certain areas of the data. While a data-aware 2D implementation will not cause any difficulty, there are issues associated with a data-agnostic 3D implementation using the current approach. Our 2D implementation captures attention as a grid overlay. If the user is allowed to rotate view of the visualization and thereby the position of the marks on the screen, the stored attention will no longer correspond to the correct marks, since they can be at another cell in the overlay. This is a general issue if the user is able to change the representation. How to address this issue is left to future work.
@@
Finally, the step to collaborative and educational uses for attention-aware visualization is not far. In situations where multiple users interact with the same visualization, AAVs could dynamically adjust to accommodate the focus and needs of different users, enhancing collaborative analysis and decision-making. In educational contexts, AAVs could tailor the presentation of information to the learner’s pace and points of interest, providing a more engaging and effective learning experience.